OpenStudy (anonymous):

Find an equation in standard form for the line described: Passing through the points (1,4) and (-3,4) If you could show the conversion from slope intercept to standard form in your work, that would help me understand this a lot!

OpenStudy (jdoe0001):

can you get the slope off those 2 points?

OpenStudy (anonymous):

Yeah, it's 0/-4

OpenStudy (anonymous):

Which means that the line is horizontal?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so our slope is 0, so let's use a point off those 2 available to us :) let's say we use ( -3, 4) and plug them in the slope-intercept form, gimme a sec

OpenStudy (anonymous):

Alright... then what?

OpenStudy (anonymous):

So basically it's 4=-3x0/4+b

OpenStudy (jdoe0001):

\(\bf y-y_1=m(x-x_1)\quad \textit{using } (\color{red}{-3, 4}) \quad y-(\color{red}{4})=(0)(x-(\color{red}{-3}))\) what do you get off that if you solve for "y"?

OpenStudy (anonymous):

you get y=4?

OpenStudy (jdoe0001):

 as you can see, as you suspected, it's a horizontal line

OpenStudy (anonymous):

So then how would this answer be converted into standard form? Or is y=4 already in standard form?

OpenStudy (jdoe0001):

standard form? looks in order to me already unless you mean y = mx+b so-called slope-intercept form

OpenStudy (anonymous):

no no, I mean the Ax+By=C form

OpenStudy (jdoe0001):

well, yes, y = 4 is standard polynomial form you only have 1 variable, and 1 independent term, or so-called constant thus you could just rewrite it as y-4 = 0